Initial state probabilities are(1/3,2/3). After one step P(X₁=1)=1/6+4/9=11/18. After the next, P(X₂=1)=(7/18)(1/2)+(11/18)(2/3)=65/108. For a0–1 variable expectation equals its probability of1. Thus27E[X₂]=65/4=16.25.
More Markov Chains questions
GATE 2020 ST Q52– A Markov chain on0,1,2 has P=pmatrix0&1&0\\.3&0&.7\\0&1&0pmatrix. Find nP(X 2n=2 X 0=2)…
GATE 2020 ST Q55– A Markov chain on0,1,2,3 has P=pmatrix1/2&1/2&0&0\\1/2&1/4&1/4&0\\0&0&1/4&3/4\\0&0&0&1pmat…
GATE 2026 ST Q18– A homogeneous Markov chain on 1,…,7 has transition matrix…
GATE 2025 ST Q51– A Markov chain on 1,2,3 has P=pmatrix0&1/2&1/2\\1/3&0&2/3\\2/5&3/5&0pmatrix and…
GATE 2024 ST Q53– A chain on 0,1,2 has P=pmatrix1/2&1/2&0\\1/2&1/2&0\\1/3&1/3&1/3pmatrix. Which hold?
GATE 2024 ST Q60– A chain on 0,1,2 has P=pmatrix0&1/2&1/2\\1/2&1/4&1/4\\1/2&1/4&1/4pmatrix and initial…